OpenStudy (anonymous):

trying to figure out what the period is for y=sec 4x?

OpenStudy (zepp):

The general formula to find the period is \[\frac{\text{Regular period of the function}}{B}\]B being y=ATrigFunction(Bx)

OpenStudy (zepp):

Can you identify the Regular period of the function and the B value of \(y=\sec(4x)\)?

OpenStudy (anonymous):

well is the regular period just pie?

OpenStudy (zepp):

The regular period for sin(x), cos(x), sec(x) and csc(x) and their inverse functions are \(2\pi\) and tan(x) and cot(x) is \(\pi\). You can use WolframAlpha to observe the core functions' periods and their many nice properties.

OpenStudy (anonymous):

awsome i really appreciate the help